A corporate bond with a face value of $100,000 was issued six years ago and there are nine years remaining until maturity. The bond pays semi-annual coupon payments of $4500, the coupon rate is 9% p.a. paid twice yearly and rates in the marketplace are 8% p.a. compounded semi-annually. What is the value of the bond today?
Answer #1

Bond Valuation: The value of bond is the present value of the expected cashflows from the bond,discounted at Yield to Maturity(YTM).

Prima facie, the bond will trade at Premium as YTM<coupon rate

Year Cash flow PVAF/PVF@4% Present Value (Cashflow*PVAF/PVF)
1-18            4,500 12.6593*                           56,966.84
18        100,000 0.4936**                           49,362.81

Current Market Price of Bonds = \sumCashflow*PVAF/PVF

= 56966.84+49362.81

= $106,329.65

Note : Since the bond makes semiannual interest payments, total no. of period is 18 (9*2), and cashflows are discounted at 4% (8/2)

*PVAF = (1-(1+r)^-n)/r

**PVF = 1 / (1+r)^n